OpenStudy (kainui):

Technique I'm trying to work out for finding the general term for recurrence relations of a special form.

OpenStudy (kainui):

\[a_{n+1}+p_n a_n = q_n\] You're given \(p_n\) and \(q_n\) as functions of n, they could be like \(n^3\) for instance. And we're also given the terms to do recurrence, \(a_0\) and \(a_1\). So what's \(a_n\)?

OpenStudy (kainui):

Here's my idea, to solve it like a differential equation, I'l make up my own series of terms, and call them \(\mu_n\) and multiply it by this recurrence relation: \[a_{n+1}+p_n a_n = q_n\] Multiply through by \(\mu_{n+1}\) \[\mu_{n+1}a_{n+1}+\mu_{n+1}p_n a_n = \mu_{n+1}q_n\] Since \(\mu_n\) are arbitrarily made up by me, I will say it satisfies this relationship: \[\mu_{n+1} p_n = -\mu_n\] Plug it in: \[\mu_{n+1}a_{n+1}-\mu_n a_n = \mu_{n+1}q_n\] So now \(\mu_{n+1}q_n\) is the difference between two consecutive terms of \(\mu_na_n\) so if we add up both sides by a bunch of consecutive terms we'll have something like this: \[\sum \mu_{n+1}a_{n+1}-\mu_n a_n =\sum \mu_{n+1}q_n\] So the left will end up collapsing as a telescoping series: \[\mu_n a_n = \sum \mu_{n+1}q_n\] Rearrange to get: \[ a_n = \frac{1}{\mu_n}\sum \mu_{n+1}q_n\] So now it looks like we have the nth term but we still have to solve for \[\mu_{n+1} p_n = -\mu_n\] and we could also plug this into that sum to make it look a little different but still the same answer: \[ a_n = \frac{-1}{\mu_n}\sum \frac{q_n}{p_n} \mu_n\] So I don't know if this is exactly the same thing as solved, but this looks pretty useful.

OpenStudy (kainui):

Yeah I'll admit some of the reasoning is kinda weird, not too sure about this telescoping series part, but anyways I guess it's simple enough to just sorta try to test it out maybe?

OpenStudy (freckles):

\[\sum \mu_{n+1}a_{n+1}-\mu_n a_n =\sum \mu_{n+1}q_n \\ \text{ wouldn't this be } \\ -\mu_o a_o=\sum \mu_{n+1} q_n\]

OpenStudy (freckles):

that seems weird though

OpenStudy (freckles):

\[\mu _n a_n-\mu_0 a_0=\sum_{k=0}^n u_{k+1} q_k\]

OpenStudy (kainui):

\[\sum \mu_{n+1}a_{n+1}-\mu_n a_n =\sum \mu_{n+1}q_n\] I'll clarify this to really mean I am looking at this: \[\sum_{k=r}^{n-1} [\mu_{k+1}a_{k+1}-\mu_k a_k] =\sum_{k=r}^{n-1} \mu_{k+1}q_k\] This left hand side will specifically collapse down to this: \[\sum_{k=r}^{n-1} [\mu_{k+1}a_{k+1}-\mu_k a_k] = \mu_na_n-\mu_ra_r\] But I haven't specified what \(r\) is so I can set it to be anything, so I guess I'm assuming there's some r I can pick that will make either \(\mu_r\) or \(a_r\) equal to 0, and that will be very simple to do if I'm given \(a_0=0\) for example. So now I have: \[a_n = \frac{1}{\mu_n}\sum_{k=r}^{n-1} \mu_{k+1}q_k\] This sum now starts at r=0 or whatever we like, but it doesn't matter because we'll end up just having to put a constant with it, cause this is the sum of differences... Kinda confusing I think hahaha. \[a_n = \frac{1}{\mu_n}\left( C + \sum_{k=r}^{n-1} \mu_{k+1}q_k \right)\] If you try to go backwards with this, and undo the summation then the C term will go away. I can explain more it's sorta weird and I've been spending some time getting comfortable with this, it's sorta like realizing that in calculus when the write: \[x^2 = \int 2x dx\] The x on the left is not the x on the right. Really it's more like: \[x^2 = \int_0^x 2t dt\] So that's why I originally wrote the summation without bounds.

OpenStudy (kainui):

Actually the way I described that was bad, we don't have to choose r=0 or anything like that. We can still have \(\mu_r a_r \ne 0\), this thing really just amounts to being our constant of 'integration'.

OpenStudy (freckles):

seems interesting

OpenStudy (freckles):

it works almost just like solving a first order linear differential equation

OpenStudy (freckles):

solving that one equation in a general way might be hard though I'm talking about the \[\mu_n=-\mu_{n+1}p_n\]

OpenStudy (kainui):

Yeah, I was thinking about it, and I think I figured out a way to solve it: \[\mu_{n+1} = \tfrac{-1}{p_n} \mu_n\] Let's just call the first index of \(p_n\) to be \(n=0\) although it doesn't have to be I guess this make it easier to follow the argument. Then we have: \[\mu_1 = \tfrac{-1}{p_0}\mu_0\] \[\mu_2 = \tfrac{-1}{p_1}\mu_1= \tfrac{-1}{p_1} \tfrac{-1}{p_0}\mu_0\] So it looks like a general form is going to be (as much as I can find without actually knowing \(p_n\): \[\mu_n =\mu_0 \frac{(-1)^n}{\prod_{k<n} p_k}\] And similarly to the arbitrariness of our choice of \(\mu\) in solving a differential equations, it seems we can define \(\mu_0=1\) unless there's some other choice that naturally seems to go well with the product of \(p_k\)s such as \(\mu_0=-1\) or something.

OpenStudy (freckles):

jenkins are whatever felma or velma or whatever her name says that seems to work I tried it for p_n=e^n... \[\mu_n=\mu_0 \cdot \frac{(-1)^n}{e^1 \cdot e^2 \cdot e^3 \cdots e^{n-1}}=\mu_0 \frac{(-1)^{n}}{e^{\frac{n(n-1)}{2}}}\] and looky http://www.wolframalpha.com/input/?i=e%5En*f%28n%2B1%29%2Bf%28n%29%3D0 that is kind of awesome

OpenStudy (freckles):

you are brilliant

OpenStudy (kainui):

Also I should note, it looks like every recurrence relation of this form can be turned into a difference equation by a pretty straight forward substitution: \[f(x+1)+p(x)f(x)=q(x)\] substitute in: \[p(x)=r(x)-1\] \[f(x+1)-f(x)+r(x)f(x)=q(x)\]\[\Delta f(x) + r(x)f(x)=q(x)\] Which is really how I thought of all this stuff to start with but it's exactly equivalent, the only difference is the shift by 1, not a big deal when talking about arbitrary functions I think but might be nice to see for fun.

OpenStudy (kainui):

HAhaha thanks xD also I like this I didn't even think about that that's a good idea to use that products turn expontents into addition thanks, I think I'll steal that and use that to pay around haha

OpenStudy (freckles):

lol

OpenStudy (freckles):

it would be nice if it paid too

OpenStudy (kainui):

hahaha right? I need a job actually xD I do too much math and not enough job searching lol

OpenStudy (freckles):

what about the nsa i think they do really fun math there not sure never worked there

OpenStudy (freckles):

I just believe they do for some odd reason

OpenStudy (kainui):

Yeah I had heard that as well actually I don't know why but it's probably true haha

OpenStudy (freckles):

i think they have to kill you or employee you if they tell you

OpenStudy (freckles):

very top secret stuff

OpenStudy (kainui):

Yeah maybe I should apply there, although I don't have a phd or masters so I don't know if they'd hire me lol

OpenStudy (freckles):

what that's crazy you know super fancy math how are you not at least a master degree person

OpenStudy (kainui):

I have a bachelor of arts in chemistry and a minor in math, I guess not having a BS gave me more free time to do whatever I wanted haha.

OpenStudy (freckles):

i'm envious

OpenStudy (kainui):

Haha well hey at least we're doing math right now and working together, this is good!

OpenStudy (freckles):

well i'm pretty sure you just solved that whole thing and I just watched I didn't do anything special

OpenStudy (freckles):

so you kinda worked alone

OpenStudy (freckles):

anyways i'm going to go hula hoop

OpenStudy (freckles):

cool problem

OpenStudy (kainui):

Well maybe to an extent but like I had already worked through most of this before I got here, I was more or less coming on here to see if other people were interested and now I basically got you up to the same level of understanding of this thing as me. The only thing I was thinking of for that product was to use factorials which are kinda boring, throwing that exponential in like that has helped a lot, plus if you think of any ideas throw 'em at me and I'll try helping you out with some stuff!

OpenStudy (freckles):

Oh you know what that one thing sorta reminded me of a geometric sequence since it can be written as \[\frac{v_{n+1}}{v_n}=\frac{-1}{p_n} \\ \text{ this reminded me of separation by variables } \\ \text{ which is everyone favorite way \to solve that one part } \\ \text{ and deriving the formula for solving first order linear differential equations } \\ \text{ you know the } v'=pv \text{ thing } \\ \text{ anyways I realize that } \prod_{k=0}^{n-1} \frac{v_{n+1}}{v_n}=\frac{v_n}{v_0} \\ \text{ so we just have \to take the product of both sides } \\ \text{ once we separate } \\ \prod_{k=0}^{n-1} \frac{v_{k+1}}{v_k}=\prod_{k=0}^{n-1} \frac{-1}{p_k} \\ \frac{v_n}{v_0}= \frac{(-1)^{n-1}}{\prod_{k=0}^{n-1} p_k} \\ v_n= v_0 \frac{(-1)^{n-1}}{\prod_{k=0}^{n-1}p_k}\]

OpenStudy (freckles):

v was easier to write then mu since mu requires two letters :p

OpenStudy (freckles):

you know instead of trying to inspect a pattern

OpenStudy (kainui):

Oh woah this is much cleaner I really like this a lot better than my sorta hand wavy guess the pattern method. Something I just realized is we have this really cool thing happening connecting sums and products that I think you already used earlier in the post to test out the product with that e^n thing. \[\ln \left( \prod a_n \right) = \sum \ln(a_n)\]

OpenStudy (freckles):

yep. i think that is actually the way wolfram came up with that one answer in that link I posted above... like they had for some reason put (-1)*constant instead of just constant I was wondering why they would do that

OpenStudy (freckles):

and yes I totally see the law of logs in action there

OpenStudy (kainui):

That negative thing was bothering me so I just changed the p(x) part to be -p(x) so that we could completely remove it and it ends up being much cleaner. \[a_{n+1}-p_n a_n = q_n\] I sorta typed this up but it's not complete yet since I had some other idea to extend this to the more general form where instead of the left-most term being \(a_{n+1}\) I instead noticed it wouldn't be too much harder to deal with it being a variable step higher, \(a_{n+k}\) and then that would admit k different solutions from k different starting points. https://www.overleaf.com/read/gkchtfggqqbn

OpenStudy (kainui):

Errr... That's actually not anything new I think it's pretty much just the same thing we already did here and nothing new I thought I had put some stuff in there but I haven't sorry

OpenStudy (freckles):

oh you are writing a paper ?

OpenStudy (freckles):

it looks pretty did you type that using winedt?